# Optimizing Optimizers for Fast Gradient-Based Learning

### v1

@@ Appendix B Instantiation of optimal optimizers
In Section 3 of the main manuscript, we introduced a practical instantiation of the theoretically derived optimal conditions for the SGD+Momentum and Adam (kingma2015adam; loshchilov2017decoupled) families of optimizers. This section gives the instantiation of the algorithms for these optimal conditions. Practical considerations are also suggested.
@@ B.1 General meta-algorithm for optimal optimizer by selection
where $a(\beta_{1},\beta_{2})$ is a scalar normalization factor, $u_{\beta_{1},\beta_{2}}=m_{\beta_{1}}/v_{\beta_{2}}^{1/2}$ is the Adam update and $v_{\beta_{2}}$ is the running average second-moment operator with parameter $\beta_{2}$ defined as $v_{\beta_{2}}[n]=\beta_{2}v_{\beta_{2}}[n-1]+(1-\beta_{2})g_{j}^{2}[n]$. During optimization, the distributions of gradients $g$, the momenta $m_{\beta}$, and the parameter updates $u$ are all time-varying, and the exact value of the expectation $\mathbb{E}$ in the equations 24 and 25 are unknown. Therefore, we approximate this by an immediate dot product between the instantaneous gradient $g[n]$ and the instantaneous parameter update $m[n]$ or $u[n]$.
This leads to Algorithm 1. The lines highlighted in blue are the ones that are different from the standard training loop. In all our experiments, we use maximal of $K=2$ candidate optimizers. In practice, this only adds less than 5% computational overhead as elaborated in Table 17.
@@
1:Candidate optimizers $\{Q_{1},Q_{2},\ldots,Q_{K}\}$ with hyperparameters $\{\beta_{1},\beta_{2},\ldots,\beta_{K}\}$
2:Initialize model and all candidate optimizers.
3:for each training step $n$ do
4:  Compute forward pass: ${\mathcal{L}}(\theta[n])$.
5:  Compute current gradient: $g[n]=-\nabla_{\theta}{\mathcal{L}}(\theta[n])$.
6:  for each candidate optimizer $Q_{k}$ with hyperparameter $\beta_{k}$ do
7:   Update internal state: $Q_{k}\leftarrow Q_{k}(g[n])$.
8:   Evaluate objective function: $J(\beta_{k},g[n])$.
9:  end for
10:  Get the optimal optimizer: $Q^{\star}=Q_{k}$ where $k=\arg\max_{k}J(\beta_{k},g[n])$.
11:  Update parameters: $\theta[n+1]\leftarrow\theta[n]-Q^{\star}g[n]$.
12:  Apply hysteresis reset for optimizer internal states to stabilize selection.
13:end for
@@
1:Learning rate $\eta$, number of candidate optimizers $K$, candidate optimizers $\{Q_{1},Q_{2},\ldots,Q_{K}\}$ with hyperparameters $\{\beta_{1},\beta_{2},\ldots,\beta_{K}\}$, respectively.
2:Initialize optimizer states $\mu_{k}\leftarrow 0$, $\forall k\in\{1,2,\ldots,K\}$, hysteresis counter $H\leftarrow 0$.
3:for each training step $n$ do
4:  $g\leftarrow\nabla_{\theta}{\mathcal{L}}(\theta)$ $\triangleright$ Standard forward-backward pass
5:  for $k=1,\ldots,K$ do
6:   $\mu_{k}\leftarrow\beta_{k}\mu_{k}+g$ $\triangleright$ Update optimizer states
7:   $J_{k}\leftarrow\sqrt{1-\beta_{k}^{2}}\,g^{\top}\mu_{k}$ $\triangleright$ Objective function for $k$-th optimizer
8:  end for
9:  $k^{\star}\leftarrow\arg\max_{k}J_{k}$ $\triangleright$ Dynamic selection of optimizer
10:  $\theta\leftarrow\theta-\eta\mu_{k^{\star}}$ $\triangleright$ Parameter update with $k^{\star}$-th optimizer
11:  if $J_{k^{\star}}<0$ for consecutive 5 steps then $\triangleright$ Hysteresis reset (stabilization trick)
12:   for $k=1,\ldots,K$ do
13:     $\mu_{k}\leftarrow 0.5\mu_{k}$
14:   end for
15:  end if
16:end for
17:return $\theta$
@@
1:Learning rate $\eta$, number of candidate optimizers $K$, candidate optimizers $\{Q_{1},Q_{2},\ldots,Q_{K}\}$ with hyperparameters $\{(\beta_{1}^{(1)},\beta_{2}^{(1)}),(\beta_{1}^{(2)},\beta_{2}^{(2)}),\ldots,(\beta_{1}^{(K)},\beta_{2}^{(K)})\}$, respectively.
2:Initialize optimizer states $\mu_{k},v_{k}\leftarrow 0$, $\forall k\in\{1,2,\ldots,K\}$, hysteresis counter $H\leftarrow 0$
3:for each calibration step $t$ do
4:  $g\leftarrow\nabla_{\theta}{\mathcal{L}}(\theta)$ $\triangleright$ Standard forward-backward pass
5:  for $k=1,\ldots,K$ do
6:   $v_{k}\leftarrow\beta_{2}^{(k)}v_{k}+(1-\beta_{2}^{(k)})g^{2}$ $\triangleright$ Update optimizer states
7:   $\mu_{k}\leftarrow\beta_{1}^{(k)}\mu_{k}+(1-\beta_{1}^{(k)})g$
8:   $c_{k}\leftarrow\sqrt{v_{k}}$ $\triangleright$ Coordinate-wise costs
9:   $u_{k}\leftarrow\mu_{k}/(c_{k}+\epsilon)$ $\triangleright$ Adam update
10:   $a_{k}\leftarrow\sqrt{1+\beta_{1}^{(k)}}/\sqrt{(1-\beta_{1}^{(k)})\sum_{j}(1/c_{k})}$ $\triangleright$ Normalization factor
11:   $J_{k}\leftarrow a_{k}\,g^{\top}u_{k}$ $\triangleright$ Objective function for $k$-th optimizer
12:  end for
13:  $k^{\star}\leftarrow\arg\max_{k}J_{k}$ $\triangleright$ Dynamic selection of optimizer
14:  $\theta\leftarrow\theta-\eta u_{k^{\star}}$ $\triangleright$ Parameter update with $k^{\star}$-th optimizer
15:  if $J_{k^{\star}}<0$ for consecutive 5 steps then $\triangleright$ Hysteresis reset (stabilization trick)
16:   for $k=1,\ldots,K$ do
17:     $\mu_{k}\leftarrow 0.5\mu_{k}$
18:   end for
19:  end if
20:end for
21:return $\theta$
@@ B.2 Practical concerns
Overall, our approach employs dynamic tuning of optimizer states. This specific implementation is similar to Grokfast (lee2024grokfast), where optimizers are viewed as filters and both gradients and parameter updates are treated as signals. We extend this idea by dynamically modulating the states of the chosen optimizer based on the instantaneous gradients. This allows us to efficiently instantiate the theoretically derived optimal conditions for existing optimizer families with minimal additional computational overhead, as demonstrated in Table 17 in Appendix C.3.
@@ B.3 Optimal SGD+Momentum and Adam instantiations
To gain further insights on the actual instantiation of this meta-algorithm for optimal optimizer selection, we provide the following two algorithms for the implementation of the optimal SGD+Momentum and Adam. Note how the hysteresis reset is applied in the last lines of the algorithms. We did not use other practical stabilization techniques that are not mentioned in the main manuscript. The code for this implementation is available at https://github.com/ironjr/optimizing-optimizers.
@@ Appendix C Implementation details and more experiments
This section provides implementation details and full results of the experiments conducted to validate the theory in Section 3. First, we provide the hyperparameters and settings for the experiments in Table 8, Table 9, Table 10, Table 11, and Table 12 of Appendix C.1. Additional experimental results are provided in Table 13, Table 14, Table 15, Table 16 of Appendix C.2. Our automatic hyperparameter tuning shows comparable performance across all datasets and models, including conventional residual networks (he2016deep), vision transformers (dosovitskiy2021image), and modern large language models (gemma2024open; llama32024herd) with or without using parameter-efficient fine-tuning methods like low-rank adaptation (LoRA) (hu2022lora). This demonstrates the practical usefulness of our framework.
@@ C.1 Implementation details
For ResNet-18 (he2016deep) on CIFAR-100 (krizhevsky2009learning) experiments, we follow the standard settings of he2016deep: 300 epochs with a learning rate decay of 0.1 at epochs 60, 120, and 160. All hyperparameters other than momentum are held fixed. We use a weight decay of $5\times 10^{-4}$, batch size of $128$, and a base learning rate of $0.1$ for SGD and $0.01$ for Adam (kingma2015adam). For our optimal Adam-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.8$ and $0.99$, to ensure that these endpoints enclose the typical range of $\beta_{1}$ values used in practice. For our optimal SGD+Momentum-type optimizers using the two-option switch, we use momentum endpoints of $0.01$ and $0.99$, again to ensure that the endpoints enclose the typical range of momentum values in practice. For optimal SGD+Momentum-type optimizers using the five-option switch, we use momentum endpoints of $0.9$, $0.95$, $0.98$, $0.99$, and $0.995$, to demonstrate the effectiveness of fine-grained control in dynamic hyperparameter tuning. In the main manuscript, we show only the results of the two-option switch for SGD+Momentum and Adam, since these do not exceed 10% of the computation time of the baseline. Our five-option switch for SGD+Momentum demonstrates that we can achieve significantly better performance than the baseline optimizer with fixed hyperparameters ($77.57\%\rightarrow 78.33\%$ test accuracy). These extended results are summarized in Tables 13 and 14 in the next section. However, current implementation of multi-option switch larger than two requires a significant amount of computation time (more than +30% compared to the baseline) and memory usage. Therefore, we did not include the results in the main manuscript, opening up a future direction for more efficient implementation.
Table 8: Hyperparameters and settings for math finetuning experiments on Gemma-2B (gemma2024open) with LoRA (hu2022lora). Values reflect the experimental script.
@@
Optimizer AdamW (kingma2015adam; loshchilov2017decoupled)
@@
Adapter type LoRA (hu2022lora)
@@
Table 9: Hyperparameters and settings for math finetuning experiments on Llama-3-8B (llama32024herd) with LoRA (hu2022lora). Values reflect the experimental script.
@@
Optimizer AdamW (kingma2015adam; loshchilov2017decoupled)
@@
Adapter type LoRA (hu2022lora)
@@
We summarize the hyperparameters and training settings for these experiments in Table 4 and Table 4 in Section 3 of the main manuscript. We use low-rank adaptation (LoRA) (hu2022lora) with a rank of $32$ and a scaling factor of $4$ for both Gemma-2B (gemma2024open) and Llama-3-8B (llama32024herd). We truncate the MetaMathQA-395K (yu2024metamath) training dataset to 100,000 examples for both models. The reported test accuracy is based on a separate, validation-only dataset, GSM8K (cobbe2021gsm8k). Additional experimental results are provided in Table 15, where we show results for baseline optimizers with different $\beta_{1}$ values we have tested. In the main manuscript, only the best baseline optimizer results are shown for brevity: $\beta_{1}=0.5$ for Gemma-2B and $\beta_{1}=0.9$ for Llama-3-8B. For our optimal AdamW-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.8$ and $0.99$, to ensure that these endpoints enclose the typical range of $\beta_{1}$ values used in practice.
Table 10: Hyperparameters and settings for the main commonsense finetuning experiments on Gemma-2B (gemma2024open) with LoRA (hu2022lora).
@@
Dataset Commonsense-170K (hu2023llmadapters)
Models tested Gemma-2B (gemma2024open)
@@
Optimizer AdamW (kingma2015adam; loshchilov2017decoupled)
@@
Adapter type LoRA (hu2022lora)
@@
Table 11: Hyperparameters and settings for the main commonsense finetuning experiments on Gemma-2B (gemma2024open) with full fine-tuning.
@@
Dataset Commonsense-170K (hu2023llmadapters)
Models tested Gemma-2B (gemma2024open)
@@
Optimizer AdamW (kingma2015adam; loshchilov2017decoupled)
@@
We summarize the hyperparameters and training settings for these experiments in Table 5 in Section 3 of the main manuscript. We use low-rank adaptation (LoRA) (hu2022lora) with a rank of $32$ and a scaling factor of $4$ for Gemma-2B (gemma2024open). We also demonstrate our optimizer with full fine-tuning on the Commonsense-170K (hu2023llmadapters) dataset. Additional experimental results are provided in Table 16 in the next section, where we show results for baseline optimizers with different $\beta_{1}$ values we have tested. In the main manuscript, only the best baseline optimizer results are shown for brevity: $\beta_{1}=0.95$ for LoRA and $\beta_{1}=0.5$ for full fine-tuning. For our optimal AdamW-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.1$ and $0.99$ for full fine-tuning and $\beta_{1}$ endpoints of $0.8$ and $0.95$ for LoRA, to ensure that these endpoints enclose the typical range of $\beta_{1}$ values used in each type of experiment in practice. After fitting the Commonsense-170K (hu2023llmadapters) dataset, we evaluate performance on various reasoning datasets that are commonly used in the LLM literature. These include BoolQ (clark2019boolq), PIQA (bisk2020piqa), Social IQA (sap2019socialiqa), HellaSwag (zellers2019hellaswag), WinoGrande (sakaguchi2019winogrande), and OBQA (hu2023llmadapters). We also report the average performance across all datasets.
Table 12: Common hyperparameters and settings for ViT-Base and ViT-Large (dosovitskiy2021image) finetuning experiments across various classification datasets.
@@
Models tested ViT-B / ViT-L (dosovitskiy2021image)
@@
Optimizer AdamW (kingma2015adam; loshchilov2017decoupled)
@@
Adapter type LoRA (hu2022lora)
@@
ViT-B and ViT-L on various classification datasets.
We summarize the hyperparameters and training settings for these experiments in Table 6 in Section 3 of the main manuscript. We tested four different configurations: ViT-Base (dosovitskiy2021image) with rank-$8$ LoRA, ViT-Base with rank-$32$ LoRA, ViT-Large with rank-$16$ LoRA, and ViT-Large with rank-$32$ LoRA. The same pre-trained weights are fine-tuned for various task-specific datasets, including Stanford Cars (krause20133d), CIFAR-100 (krizhevsky2009learning), CUB-200 (wah2011cub200), DTD (cimpoi2014describing), Food-101 (bossard2014food101), RESISC45 (cheng2017resisc45), and SUN397 (xiao2010sun). For our optimal AdamW-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.5$ and $0.99$ for all configurations. This covers the working range of $\beta_{1}$ values used in practice for each type of experiment. We also report the average performance across all datasets.
C.2 Extended experimental results
ResNet-18 on CIFAR-100.
Figure 3 in the main manuscript demonstrates how optimizer hyperparameters affect the final training loss and validation accuracy, and how our optimal optimizers compare to baseline optimizers with fixed hyperparameters. Here, the complete results are provided in Tables 13 and 14. Figure 4 further compares the training curves of baseline optimizers and our optimal optimizers. For clarity, each baseline plot shows only the best run, and only baselines with robust momentum values are displayed. Specifically, we show results for $\beta\in[0.8,0.99]$ for SGD+Momentum and $\beta_{1}\in[0.8,0.99]$ for Adam (kingma2015adam). Although the abrupt learning rate decay of the scheduler introduces perturbations that are not considered in our theory, our implementation of the greedy optimal optimizer generally reduces the training loss rapidly and achieves better performance than the baselines. Moreover, the complete results in Tables 13 and 14 show that by increasing the number of selectable options from two to five, our instantiation of the greedy optimal optimizers achieves significantly better performance than baseline optimizers with any fixed hyperparameters. This opens up new opportunities for research into the dynamic hyperparameter tuning framework, which is first enabled by our theory.
@@ C.2 Extended experimental results
Figure 4: Demonstration of Corollaries 5 and 6. Our instantiations of optimal optimizers are compared with baselines having fixed hyperparameters on the CIFAR-100 dataset (krizhevsky2009learning) with ResNet-18 (he2016deep), following the standard settings of he2016deep. The line and shaded area indicate the mean and standard deviation over 10 runs. For clear visualization, each baseline plot shows only the best run. For SGD+Momentum, momentum below 0.8 showed suboptimal performance.
@@
Method GSM8K acc. (%) Train loss
$\beta_{1}=0.5$ 52.57 $\pm$ 1.10 0.2080 $\pm$ 0.0004
$\beta_{1}=0.8$ 52.31 $\pm$ 1.00 0.2080 $\pm$ 0.0004
$\beta_{1}=0.9$ 51.76 $\pm$ 0.99 0.2081 $\pm$ 0.0004
$\beta_{1}=0.95$ 51.12 $\pm$ 0.77 0.2085 $\pm$ 0.0004
$\beta_{1}=0.98$ 51.25 $\pm$ 0.38 0.2093 $\pm$ 0.0005
$\beta_{1}=0.99$ 50.97 $\pm$ 0.68 0.2103 $\pm$ 0.0004
Ours 52.77 $\pm$ 0.93 0.2084 $\pm$ 0.0003

### v2

@@ Appendix B Instantiation of Optimal Optimizers
In Section 3 of the main manuscript, we introduced a practical instantiation of the theoretically derived optimal conditions for the SGD+Momentum and Adam (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) families of optimizers. This section gives the instantiation of the algorithms for these optimal conditions. Practical considerations are also suggested.
@@ B.1 General meta-algorithm for optimal optimizer by selection
where $a(\beta_{1},\beta_{2})$ is a scalar normalization factor, $u_{\beta_{1},\beta_{2}}=m_{\beta_{1}}/v_{\beta_{2}}^{1/2}$ is the Adam update and $v_{\beta_{2}}$ is the running average second-moment operator with parameter $\beta_{2}$ defined as $v_{\beta_{2}}[n]=\beta_{2}v_{\beta_{2}}[n-1]+(1-\beta_{2})g_{j}^{2}[n]$. During optimization, the distributions of gradients $g$, the momenta $m_{\beta}$, and the parameter updates $u$ are all time-varying, and the exact value of the expectation $\mathbb{E}$ in the equations 27 and 28 are unknown. Therefore, we approximate this by an immediate dot product between the instantaneous gradient $g[n]$ and the instantaneous parameter update $m[n]$ or $u[n]$.
This leads to Algorithm 1. The lines highlighted in purple are the ones that are different from the standard training loop. In all our experiments, we use maximal of $K=2$ candidate optimizers. In practice, this only adds less than 5% computational overhead as elaborated in Table 18.
@@
Input: Candidate optimizers $\{Q_{1},Q_{2},\ldots,Q_{K}\}$ with hyperparameters $\{\beta_{1},\beta_{2},\ldots,\beta_{K}\}$
 Initialize model and all candidate optimizers.
for each training step $n$ do
  Compute forward pass: ${\mathcal{L}}(\theta[n])$.
  Compute current gradient: $g[n]=-\nabla_{\theta}{\mathcal{L}}(\theta[n])$.
  for each candidate optimizer $Q_{k}$ with hyperparameter $\beta_{k}$ do
   Update internal state: $Q_{k}\leftarrow Q_{k}(g[n])$.
   Evaluate objective function: $J(\beta_{k},g[n])$.
  end for
  Get the optimal optimizer: $Q^{\star}=Q_{k}$ where $k=\arg\max_{k}J(\beta_{k},g[n])$.
  Update parameters: $\theta[n+1]\leftarrow\theta[n]-Q^{\star}g[n]$.
  Apply hysteresis reset for optimizer internal states to stabilize selection.
end for
@@
Input: Learning rate $\eta$, number of candidate optimizers $K$, candidate optimizers $\{Q_{1},Q_{2},\ldots,Q_{K}\}$ with hyperparameters $\{\beta_{1},\beta_{2},\ldots,\beta_{K}\}$, respectively.
 Initialize optimizer states $\mu_{k}\leftarrow 0$, $\forall k\in\{1,2,\ldots,K\}$, hysteresis counter $H\leftarrow 0$.
for each training step $n$ do
  $g\leftarrow\nabla_{\theta}{\mathcal{L}}(\theta)$ $\triangleright$ Standard forward-backward pass
  for $k=1,\ldots,K$ do
   $\mu_{k}\leftarrow\beta_{k}\mu_{k}+g$ $\triangleright$ Update optimizer states
   $J_{k}\leftarrow\sqrt{1-\beta_{k}^{2}}\,g^{\top}\mu_{k}$ $\triangleright$ Objective function for $k$-th optimizer
  end for
  $k^{\star}\leftarrow\arg\max_{k}J_{k}$ $\triangleright$ Dynamic selection of optimizer
  $\theta\leftarrow\theta-\eta\mu_{k^{\star}}$ $\triangleright$ Parameter update with $k^{\star}$-th optimizer
  if $J_{k^{\star}}<0$ for consecutive 5 steps then
   for $k=1,\ldots,K$ do
    $\mu_{k}\leftarrow 0.5\mu_{k}$ $\triangleright$ Hysteresis reset (stabilization)
   end for
  end if
end for
return $\theta$
@@
Input: Learning rate $\eta$, number of candidate optimizers $K$, candidate optimizers $\{Q_{1},Q_{2},\ldots,Q_{K}\}$ with hyperparameters $\{(\beta_{1}^{(1)},\beta_{2}^{(1)}),(\beta_{1}^{(2)},\beta_{2}^{(2)}),\ldots,(\beta_{1}^{(K)},\beta_{2}^{(K)})\}$, respectively.
 Initialize optimizer states $\mu_{k},v_{k}\leftarrow 0$, $\forall k\in\{1,2,\ldots,K\}$, hysteresis counter $H\leftarrow 0$
for each calibration step $t$ do
  $g\leftarrow\nabla_{\theta}{\mathcal{L}}(\theta)$ $\triangleright$ Standard forward-backward pass
  for $k=1,\ldots,K$ do
   $v_{k}\leftarrow\beta_{2}^{(k)}v_{k}+(1-\beta_{2}^{(k)})g^{2}$ $\triangleright$ Update second moment
   $\mu_{k}\leftarrow\beta_{1}^{(k)}\mu_{k}+(1-\beta_{1}^{(k)})g$ $\triangleright$ Update first moment
   $c_{k}\leftarrow\sqrt{v_{k}}$ $\triangleright$ Coordinate-wise costs
   $u_{k}\leftarrow\mu_{k}/(c_{k}+\epsilon)$ $\triangleright$ Adam update
   $a_{k}\leftarrow\sqrt{1+\beta_{1}^{(k)}}/\sqrt{(1-\beta_{1}^{(k)})\sum_{j}(1/c_{k})}$ $\triangleright$ Normalization factor
   $J_{k}\leftarrow a_{k}\,g^{\top}u_{k}$ $\triangleright$ Objective function for $k$-th optimizer
  end for
  $k^{\star}\leftarrow\arg\max_{k}J_{k}$ $\triangleright$ Dynamic selection of optimizer
  $\theta\leftarrow\theta-\eta u_{k^{\star}}$ $\triangleright$ Parameter update with $k^{\star}$-th optimizer
  if $J_{k^{\star}}<0$ for consecutive 5 steps then
   for $k=1,\ldots,K$ do
    $\mu_{k}\leftarrow 0.5\mu_{k}$ $\triangleright$ Hysteresis reset (stabilization)
   end for
  end if
end for
return $\theta$
@@ B.2 Practical concerns
Overall, our approach employs dynamic tuning of optimizer states. This specific implementation is similar to Grokfast (Lee et al., 2024), where optimizers are viewed as filters and both gradients and parameter updates are treated as signals. We extend this idea by dynamically modulating the states of the chosen optimizer based on the instantaneous gradients. This allows us to efficiently instantiate the theoretically derived optimal conditions for existing optimizer families with minimal additional computational overhead, as demonstrated in Table 18 in Appendix C.3.
@@ B.3 Optimal SGD+Momentum and Adam instantiations
To gain further insights on the actual instantiation of this meta-algorithm for optimal optimizer selection, we provide the following two algorithms for the implementation of the optimal SGD+Momentum and Adam. Note how the hysteresis reset is applied in the last lines of the algorithms. We did not use other practical stabilization techniques that are not mentioned in the main manuscript. The code for this implementation is available at https://github.com/ironjr/loot.
@@ Appendix C Implementation Details and More Experiments
This section provides implementation details and full results of the experiments conducted to validate the theory in Section 3. First, we provide the hyperparameters and settings for the experiments in Table 9, Table 10, Table 11, Table 12, and Table 13 of Appendix C.1. Additional experimental results are provided in Table 15, Table 15, Table 17, Table 16 of Appendix C.2. Our automatic hyperparameter tuning shows comparable performance across all datasets and models, including conventional residual networks (He et al., 2016), vision transformers (Dosovitskiy et al., 2021), and modern large language models (Gemma Team and others, 2023; Grattafiori and others, 2024) with or without using parameter-efficient fine-tuning methods like low-rank adaptation (LoRA) (Hu et al., 2022). This demonstrates the practical usefulness of our framework.
@@ C.1 Implementation details
For ResNet-18 (He et al., 2016) on CIFAR-100 (Krizhevsky, 2009) experiments, we follow the standard settings of (He et al., 2016): 300 epochs with a learning rate decay of 0.1 at epochs 60, 120, and 160. All hyperparameters other than momentum are held fixed. We use a weight decay of $5\times 10^{-4}$, batch size of $128$, and a base learning rate of $0.1$ for SGD and $0.01$ for Adam (Kingma and Ba, 2015). For our optimal Adam-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.8$ and $0.99$, to ensure that these endpoints enclose the typical range of $\beta_{1}$ values used in practice. For our optimal SGD+Momentum-type optimizers using the two-option switch, we use momentum endpoints of $0.01$ and $0.99$, again to ensure that the endpoints enclose the typical range of momentum values in practice. For optimal SGD+Momentum-type optimizers using the five-option switch, we use momentum endpoints of $0.9$, $0.95$, $0.98$, $0.99$, and $0.995$, to demonstrate the effectiveness of fine-grained control in dynamic hyperparameter tuning. In the main manuscript, we show only the results of the two-option switch for SGD+Momentum and Adam, since these do not exceed 10% of the computation time of the baseline. Our five-option switch for SGD+Momentum demonstrates that we can achieve significantly better performance than the baseline optimizer with fixed hyperparameters ($77.57\%\rightarrow 78.33\%$ test accuracy). These extended results are summarized in Tables 15 and 15 in the next section. However, current implementation of multi-option switch larger than two requires a significant amount of computation time (more than +30% compared to the baseline) and memory usage. Therefore, we did not include the results in the main manuscript, opening up a future direction for more efficient implementation.
Table 9: Hyperparameters and settings for math finetuning experiments on Gemma-2B (Gemma Team and others, 2023) with LoRA (Hu et al., 2022). Values reflect the experimental script.
@@
Optimizer AdamW (Kingma and Ba, 2015; Loshchilov and Hutter, 2019)
@@
Adapter type LoRA (Hu et al., 2022)
@@
Table 10: Hyperparameters and settings for math finetuning experiments on Llama-3-8B (Grattafiori and others, 2024) with LoRA (Hu et al., 2022). Values reflect the experimental script.
@@
Optimizer AdamW (Kingma and Ba, 2015; Loshchilov and Hutter, 2019)
@@
Adapter type LoRA (Hu et al., 2022)
@@
We summarize the hyperparameters and training settings for these experiments in Table 6 and Table 6 in Section 3 of the main manuscript. We use low-rank adaptation (LoRA) (Hu et al., 2022) with a rank of $32$ and a scaling factor of $4$ for both Gemma-2B (Gemma Team and others, 2023) and Llama-3-8B (Grattafiori and others, 2024). We truncate the MetaMathQA-395K (Yu et al., 2024a) training dataset to 100,000 examples for both models. The reported test accuracy is based on a separate, validation-only dataset, GSM8K (Cobbe et al., 2021). Additional experimental results are provided in Table 17, where we show results for baseline optimizers with different $\beta_{1}$ values we have tested. In the main manuscript, only the best baseline optimizer results are shown for brevity: $\beta_{1}=0.5$ for Gemma-2B and $\beta_{1}=0.9$ for Llama-3-8B. For our optimal AdamW-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.8$ and $0.99$, to ensure that these endpoints enclose the typical range of $\beta_{1}$ values used in practice.
Table 11: Hyperparameters and settings for the main commonsense finetuning experiments on Gemma-2B (Gemma Team and others, 2023) with LoRA (Hu et al., 2022).
@@
Dataset Commonsense-170K (Hu et al., 2023)
Models tested Gemma-2B (Gemma Team and others, 2023)
@@
Optimizer AdamW (Kingma and Ba, 2015; Loshchilov and Hutter, 2019)
@@
Adapter type LoRA (Hu et al., 2022)
@@
Table 12: Hyperparameters and settings for the main commonsense finetuning experiments on Gemma-2B (Gemma Team and others, 2023) with full fine-tuning.
@@
Dataset Commonsense-170K (Hu et al., 2023)
Models tested Gemma-2B (Gemma Team and others, 2023)
@@
Optimizer AdamW (Kingma and Ba, 2015; Loshchilov and Hutter, 2019)
@@
We summarize the hyperparameters and training settings for these experiments in Table 3 in Section 3 of the main manuscript. We use low-rank adaptation (LoRA) (Hu et al., 2022) with a rank of $32$ and a scaling factor of $4$ for Gemma-2B (Gemma Team and others, 2023). We also demonstrate our optimizer with full fine-tuning on the Commonsense-170K (Hu et al., 2023) dataset. Additional experimental results are provided in Table 16 in the next section, where we show results for baseline optimizers with different $\beta_{1}$ values we have tested. In the main manuscript, only the best baseline optimizer results are shown for brevity: $\beta_{1}=0.95$ for LoRA and $\beta_{1}=0.5$ for full fine-tuning. For our optimal AdamW-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.1$ and $0.99$ for full fine-tuning and $\beta_{1}$ endpoints of $0.8$ and $0.95$ for LoRA, to ensure that these endpoints enclose the typical range of $\beta_{1}$ values used in each type of experiment in practice. After fitting the Commonsense-170K (Hu et al., 2023) dataset, we evaluate performance on various reasoning datasets that are commonly used in the LLM literature. These include BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), Social IQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), and OBQA (Hu et al., 2023). We also report the average performance across all datasets.
Table 13: Common hyperparameters and settings for ViT-Base and ViT-Large (Dosovitskiy et al., 2021) finetuning experiments across various classification datasets.
@@
Models tested ViT-B / ViT-L (Dosovitskiy et al., 2021)
@@
Optimizer AdamW (Kingma and Ba, 2015; Loshchilov and Hutter, 2019)
@@
Adapter type LoRA (Hu et al., 2022)
@@
Figure 4: Demonstration of Corollaries 3.3 and 3.4. Our instantiations of optimal optimizers are compared with baselines having fixed hyperparameters on the CIFAR-100 dataset (Krizhevsky, 2009) with ResNet-18 (He et al., 2016), following the standard settings of (He et al., 2016). The line and shaded area indicate the mean and standard deviation over 10 runs. For clear visualization, each baseline plot shows only the best run. For SGD+Momentum, momentum below 0.8 showed suboptimal performance.
@@
ViT-B and ViT-L on various classification datasets.
We summarize the hyperparameters and training settings for these experiments in Table 4 in Section 3 of the main manuscript. We tested four different configurations: ViT-Base (Dosovitskiy et al., 2021) with rank-$8$ LoRA, ViT-Base with rank-$32$ LoRA, ViT-Large with rank-$16$ LoRA, and ViT-Large with rank-$32$ LoRA. The same pre-trained weights are fine-tuned for various task-specific datasets, including Stanford Cars (Krause et al., 2013), CIFAR-100 (Krizhevsky, 2009), CUB-200 (Wah et al., 2011), DTD (Cimpoi et al., 2014), Food-101 (Bossard et al., 2014), RESISC45 (Cheng et al., 2017), and SUN397 (Xiao et al., 2010). For our optimal AdamW-type optimizers using the two-option switch, we use $\beta_{1}$ endpoints of $0.5$ and $0.99$ for all configurations. This covers the working range of $\beta_{1}$ values used in practice for each type of experiment. We also report the average performance across all datasets.
C.2 Extended experimental results
ResNet-18 on CIFAR-100.
Figure 3 in the main manuscript demonstrates how optimizer hyperparameters affect the final training loss and validation accuracy, and how our optimal optimizers compare to baseline optimizers with fixed hyperparameters. Here, the complete results are provided in Tables 15 and 15. Figure 4 further compares the training curves of baseline optimizers and our optimal optimizers. For clarity, each baseline plot shows only the best run, and only baselines with robust momentum values are displayed. Specifically, we show results for $\beta\in[0.8,0.99]$ for SGD+Momentum and $\beta_{1}\in[0.8,0.99]$ for Adam (Kingma and Ba, 2015). Although the abrupt learning rate decay of the scheduler introduces perturbations that are not considered in our theory, our implementation of the greedy optimal optimizer generally reduces the training loss rapidly and achieves better performance than the baselines. Moreover, the complete results in Tables 15 and 15 show that by increasing the number of selectable options from two to five, our instantiation of the greedy optimal optimizers achieves significantly better performance than baseline optimizers with any fixed hyperparameters. This opens up new opportunities for research into the dynamic hyperparameter tuning framework, which is first enabled by our theory.